OpenStudy (anonymous):

5/8x+1/16x=5/16+x the solution is x=? help please??

OpenStudy (anonymous):

is the x in the denominator?

OpenStudy (angela210793):

5/8x+1/16x=5/16+x (10+1)/16x=5/(16+x) 11(16+x)=5*16x 176+11x=80x 176=69x x=176/69

OpenStudy (anonymous):

5 1 5 --- + ----- = ----- + x 8x 16x 16 Multiply all terms by 16x, we get 10 + 1 = 5x + 16x^2 11 = 5x + 16x^2 16x^2 + 5x - 11 = 0 16x^2 + 16x - 11x -11 = 0 16x(x + 1) -11(x + 1) = 0 (16x - 11)(x + 1) = 0 This is possible if either (16x - 11) or (x + 1) is zero If 16x - 11 = 0 ==> 16x = 11 ==> x = 11/16 If x + 1 = 0 ==> x = -1 Therefore x = 11/16, -1

OpenStudy (anonymous):

harikat so 176/69 is wrong

OpenStudy (angela210793):

is x in the denom or not? (I've solved as it is)

OpenStudy (anonymous):

u hv to confirm if I hv written the question correct at the beginning or angela... Who has written your question properly???

OpenStudy (anonymous):

angela

OpenStudy (angela210793):

Next time it would be better if u used some brackets :):)

OpenStudy (anonymous):

ok thanks sorry.. a little rusty with this algebra

OpenStudy (angela210793):

:)